OpenStudy (anonymous):

is this equation true or false? \[28-4\sqrt{2}=24\sqrt{2}\]. explaine why.

OpenStudy (anonymous):

this is false 28√2 - 4 √2 = 24√2

OpenStudy (anonymous):

wrong, because 28 is a different term than \[4\sqrt{2}\] therefore they can't be combined. the simplified answer would be the 28-4 sqrt 2

OpenStudy (anonymous):

you can treat radicals like letters : 28x - 4x = 24x